# IR2: Information Regularization for Information Retrieval thanks:   Accepted by LREC-COLING 2024

### v1

@@ 5.1.   Results on WhatsThatBook
5.2.   Results on DORIS-MAE
E5-Large-v2 RoBERTa
Method
R@5
R@20
NDCG@10
MRR@10
R@5
R@20
NDCG@10
MRR@10
Pretrained
14.67
42.15
71.98
14.34
11.96
34.77
66.86
8.56
Contrastive Fine-tuning
15.28
39.87
72.51
13.43
14.75
43.32
73.55
20.79
Promptagator
16.18
45.59
73.95
14.83
13.03
42.88
72.23
16.82
Doc40% reg
14.90 $\downarrow$ -1.28
46.55 $\uparrow$ 0.96
74.28 $\uparrow$ 0.33
14.87
12.54
44.89 $\uparrow$ 2.01
72.94 $\uparrow$ 0.71
18.16 $\uparrow$ 1.34
Queryreg $\circ$ Doc40% reg
15.85
45.67
73.96
16.10 $\uparrow$ 1.27
14.66 $\uparrow$ 1.63
47.41 $\uparrow$ 4.53
74.04 $\uparrow$ 1.81
18.81 $\uparrow$ 1.99
Doc60% reg
15.23 $\downarrow$ -0.95
46.57 $\uparrow$ 0.98
74.98 $\uparrow$ 1.03
15.14
14.11 $\uparrow$ 1.08
43.99 $\uparrow$ 1.11
72.94 $\uparrow$ 0.71
16.84
Queryreg $\circ$ Doc60% reg
15.31 $\downarrow$ -0.87
48.16 $\uparrow$ 2.57
74.90 $\uparrow$ 0.95
15.43 $\uparrow$ 0.60
14.75 $\uparrow$ 1.72
47.25 $\uparrow$ 4.37
74.61 $\uparrow$ 2.38
20.48 $\uparrow$ 3.66
Doc80% reg
14.68 $\downarrow$ -1.50
43.89 $\downarrow$ -1.70
74.13 $\uparrow$ 0.18
13.51 $\downarrow$ -1.32
14.47 $\uparrow$ 1.44
42.47
72.61 $\uparrow$ 0.38
15.10 $\downarrow$ -1.72
Queryreg $\circ$ Doc80% reg
15.63 $\downarrow$ -0.55
46.75 $\uparrow$ 1.16
75.08 $\uparrow$ 1.13
16.62 $\uparrow$ 1.79
15.84 $\uparrow$ 2.81
46.96 $\uparrow$ 4.08
74.71 $\uparrow$ 2.48
20.49 $\uparrow$ 3.67
Instrreg
16.38
46.73 $\uparrow$ 1.14
74.42 $\uparrow$ 0.47
17.06 $\uparrow$ 2.23
13.80 $\uparrow$ 0.77
45.78 $\uparrow$ 2.90
73.53 $\uparrow$ 1.30
18.24 $\uparrow$ 1.42
Queryreg $\circ$ Instrreg
15.76 $\downarrow$ -0.42
48.46 $\uparrow$ 2.87
76.02 $\uparrow$ 2.07
20.86 $\uparrow$ 6.03
15.20 $\uparrow$ 2.17
47.44 $\uparrow$ 4.56
74.08 $\uparrow$ 1.85
19.31 $\uparrow$ 2.49
SimCSE SPECTER-v2
Method
R@5
R@20
NDCG@10
MRR@10
R@5
R@20
NDCG@10
MRR@10
Pretrained
14.38
41.83
70.81
24.83
13.59
41.92
71.46
21.74
Contrastive Fine-tuning
15.60
44.99
73.18
22.41
14.56
45.64
71.46
20.49
Promptagator
16.35
46.69
73.97
16.54
14.13
39.22
71.55
22.37
Doc40% reg
17.70 $\uparrow$ 1.35
45.30 $\downarrow$ -1.39
74.49 $\uparrow$ 0.52
16.49
15.39 $\uparrow$ 1.26
43.39 $\uparrow$ 4.17
72.11 $\uparrow$ 0.56
21.66 $\downarrow$ -0.71
Queryreg $\circ$ Doc40% reg
16.06
45.46 $\downarrow$ -1.23
73.23 $\downarrow$ -0.74
16.56
14.19
42.66 $\uparrow$ 3.44
71.49
21.40 $\downarrow$ -0.97
Doc60% reg
17.08 $\uparrow$ 0.73
45.33 $\downarrow$ -1.36
74.70 $\uparrow$ 0.73
17.75 $\uparrow$ 1.21
15.41 $\uparrow$ 1.28
44.19 $\uparrow$ 4.97
72.81 $\uparrow$ 1.26
21.35 $\downarrow$ -1.02
Queryreg $\circ$ Doc60% reg
17.04 $\uparrow$ 0.69
46.68
74.65 $\uparrow$ 0.68
18.32 $\uparrow$ 1.78
14.95 $\uparrow$ 0.82
43.30 $\uparrow$ 4.08
72.19 $\uparrow$ 0.64
21.60 $\downarrow$ -0.77
Doc80% reg
15.01 $\downarrow$ -1.34
42.53 $\downarrow$ -4.16
73.69 $\downarrow$ -0.28
18.68 $\uparrow$ 2.14
17.02 $\uparrow$ 2.89
44.97 $\uparrow$ 5.75
73.94 $\uparrow$ 2.39
22.78 $\uparrow$ 0.41
Queryreg $\circ$ Doc80% reg
15.92
45.70 $\downarrow$ -0.99
73.86
18.54 $\uparrow$ 2.00
14.70 $\uparrow$ 0.57
44.66 $\uparrow$ 5.44
73.07 $\uparrow$ 1.52
22.65
Instrreg
16.50
46.68
75.60 $\uparrow$ 1.63
19.71 $\uparrow$ 3.17
14.73 $\uparrow$ 0.60
41.50 $\uparrow$ 2.28
73.06 $\uparrow$ 1.51
21.05 $\downarrow$ -1.32
Queryreg $\circ$ Instrreg
16.06
47.61 $\uparrow$ 0.92
75.43 $\uparrow$ 1.46
22.14 $\uparrow$ 5.60
14.50 $\uparrow$ 0.37
41.44 $\uparrow$ 2.22
72.46 $\uparrow$ 0.91
22.26
@@ 5.3.   Results on ArguAna
Method
R@5
R@20
NDCG@10
MRR@10
R@5
R@20
NDCG@10
MRR@10
Pretrained
59.37
88.16
47.75
38.96
30.51
49.68
23.56
18.49
Contrastive Fine-tuning
70.69
94.77
56.15
47.13
59.12
84.67
47.57
39.65
Promptagator
69.41
94.11
55.14
46.67
61.51
88.19
49.80
41.85
Doc40% reg
78.72 $\uparrow$ 9.31
96.46 $\uparrow$ 2.35
64.24 $\uparrow$ 9.10
56.45 $\uparrow$ 9.78
67.38 $\uparrow$ 5.87
91.03 $\uparrow$ 2.84
54.54 $\uparrow$ 4.74
46.43 $\uparrow$ 4.58
Queryreg $\circ$ Doc40% reg
77.76 $\uparrow$ 8.35
96.21 $\uparrow$ 2.10
63.00 $\uparrow$ 7.86
55.24 $\uparrow$ 8.57
67.29 $\uparrow$ 5.78
91.08 $\uparrow$ 2.89
54.04 $\uparrow$ 4.24
45.85 $\uparrow$ 4.00
Doc60% reg
78.16 $\uparrow$ 8.75
96.41 $\uparrow$ 2.30
64.41 $\uparrow$ 9.27
56.57 $\uparrow$ 9.90
67.51 $\uparrow$ 6.00
90.77 $\uparrow$ 2.58
54.85 $\uparrow$ 5.05
46.85 $\uparrow$ 5.00
Queryreg $\circ$ Doc60% reg
78.34 $\uparrow$ 8.93
96.37 $\uparrow$ 2.26
64.10 $\uparrow$ 8.96
56.24 $\uparrow$ 9.57
67.71 $\uparrow$ 6.20
90.99 $\uparrow$ 2.80
54.41 $\uparrow$ 4.61
46.34 $\uparrow$ 4.49
Doc80% reg
77.37 $\uparrow$ 7.96
95.57 $\uparrow$ 1.46
63.14 $\uparrow$ 8.00
55.16 $\uparrow$ 8.49
67.82 $\uparrow$ 6.31
90.18 $\uparrow$ 1.99
54.62 $\uparrow$ 4.82
46.79 $\uparrow$ 4.94
Queryreg $\circ$ Doc80% reg
78.44 $\uparrow$ 9.03
96.25 $\uparrow$ 2.14
64.36 $\uparrow$ 9.22
56.49 $\uparrow$ 9.82
67.38 $\uparrow$ 5.87
90.70 $\uparrow$ 2.51
54.53 $\uparrow$ 4.73
46.59 $\uparrow$ 4.74
Instrreg
72.06 $\uparrow$ 2.65
95.70 $\uparrow$ 1.59
57.76 $\uparrow$ 2.62
49.17 $\uparrow$ 2.50
64.47 $\uparrow$ 2.96
90.09 $\uparrow$ 1.90
51.49 $\uparrow$ 1.69
43.20 $\uparrow$ 1.35
Queryreg $\circ$ Instrreg
71.59 $\uparrow$ 2.18
95.18 $\uparrow$ 1.07
56.89 $\uparrow$ 1.75
48.37 $\uparrow$ 1.70
65.14 $\uparrow$ 3.63
90.26 $\uparrow$ 2.07
52.25 $\uparrow$ 2.45
43.89 $\uparrow$ 2.04
@@
Method
R@5
R@20
NDCG@10
MRR@10
R@5
R@20
NDCG@10
MRR@10
Pretrained
49.03
81.12
39.23
30.35
38.91
71.79
30.91
23.23
Contrastive Fine-tuning
36.70
64.79
28.85
22.12
41.90
75.70
33.20
25.26
Promptagator
69.49
92.54
56.38
47.97
38.64
75.15
30.24
21.77
Doc40% reg
74.92 $\uparrow$ 5.43
93.36 $\uparrow$ 0.82
60.24 $\uparrow$ 3.86
52.42 $\uparrow$ 4.45
56.09 $\uparrow$ 17.45
86.32 $\uparrow$ 11.17
45.00 $\uparrow$ 14.76
36.30 $\uparrow$ 14.53
Queryreg $\circ$ Doc40% reg
76.07 $\uparrow$ 6.58
93.97 $\uparrow$ 1.43
61.38 $\uparrow$ 5.00
53.57 $\uparrow$ 5.60
51.28 $\uparrow$ 12.64
83.33 $\uparrow$ 8.18
40.59 $\uparrow$ 10.35
31.81 $\uparrow$ 10.04
Doc60% reg
73.20 $\uparrow$ 3.71
92.10
58.46 $\uparrow$ 2.08
50.58 $\uparrow$ 2.61
58.87 $\uparrow$ 20.23
86.93 $\uparrow$ 11.78
46.55 $\uparrow$ 16.31
37.88 $\uparrow$ 16.11
Queryreg $\circ$ Doc60% reg
75.75 $\uparrow$ 6.26
93.19 $\uparrow$ 0.65
60.94 $\uparrow$ 4.56
53.24 $\uparrow$ 5.27
54.22 $\uparrow$ 15.58
84.97 $\uparrow$ 9.82
42.90 $\uparrow$ 12.66
34.21 $\uparrow$ 12.44
Doc80% reg
70.38
89.67 $\downarrow$ -2.87
55.76
47.97
61.02 $\uparrow$ 22.38
87.48 $\uparrow$ 12.33
48.68 $\uparrow$ 18.44
40.24 $\uparrow$ 18.47
Queryreg $\circ$ Doc80% reg
74.10 $\uparrow$ 4.61
92.59
59.79 $\uparrow$ 3.41
52.18 $\uparrow$ 4.21
57.55 $\uparrow$ 18.91
86.65 $\uparrow$ 11.50
46.01 $\uparrow$ 15.77
37.44 $\uparrow$ 15.67
Instrreg
70.97
93.91 $\uparrow$ 1.37
57.41
48.80
43.26 $\uparrow$ 4.62
80.16 $\uparrow$ 5.01
34.80 $\uparrow$ 4.56
26.07 $\uparrow$ 4.30
Queryreg $\circ$ Instrreg
71.46 $\uparrow$ 1.97
93.78 $\uparrow$ 1.24
58.16 $\uparrow$ 1.78
49.78 $\uparrow$ 1.81
38.57
77.12 $\uparrow$ 1.97
31.72 $\uparrow$ 1.48
23.15 $\uparrow$ 1.38
@@ 6.   Limitations
Expanding on the current research, there is an opportunity to explore information regularization in a more granular context, possibly by introducing more sophisticated measures than the existing $p\%$ parameter. Understanding the synergy between different regularization strategies could also unveil new insights into optimal model training practices.
Our experiments relied on models which were accessed through API. We use stable checkpoints for these models, gpt-4-0613 and gpt-3.5-turbo-0301, which will allow for reproducibility as long as these checkpoints are maintained.
7.   Bibliographical References
[…]
  • Bonifacio et al. (2022) Luiz Bonifacio, Hugo Abonizio, Marzieh Fadaee, and Rodrigo Nogueira. 2022. Inpars: Unsupervised dataset generation for information retrieval. In Proceedings of the 45th International ACM SIGIR Conference on Research and Development in Information Retrieval, pages 2387–2392.
  • Chen et al. (2020) Ting Chen, Simon Kornblith, Mohammad Norouzi, and Geoffrey Hinton. 2020. A simple framework for contrastive learning of visual representations. In International conference on machine learning, pages 1597–1607. PMLR.
  • Chowdhery et al. (2022) Aakanksha Chowdhery, Sharan Narang, Jacob Devlin, Maarten Bosma, Gaurav Mishra, Adam Roberts, Paul Barham, Hyung Won Chung, Charles Sutton, Sebastian Gehrmann, et al. 2022. Palm: Scaling language modeling with pathways. arXiv preprint arXiv:2204.02311.
  • Chuang et al. (2022) Yung-Sung Chuang, Rumen Dangovski, Hongyin Luo, Yang Zhang, Shiyu Chang, Marin Soljacic, Shang-Wen Li, Scott Yih, Yoon Kim, and James Glass. 2022. DiffCSE: Difference-based contrastive learning for sentence embeddings. In Proceedings of the 2022 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, pages 4207–4218, Seattle, United States. Association for Computational Linguistics.
  • Cohan et al. (2020) Arman Cohan, Sergey Feldman, Iz Beltagy, Doug Downey, and Daniel S Weld. 2020. Specter: Document-level representation learning using citation-informed transformers. In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics, pages 2270–2282.
  • Dai et al. (2022) Zhuyun Dai, Vincent Y Zhao, Ji Ma, Yi Luan, Jianmo Ni, Jing Lu, Anton Bakalov, Kelvin Guu, Keith Hall, and Ming-Wei Chang. 2022. Promptagator: Few-shot dense retrieval from 8 examples. In The Eleventh International Conference on Learning Representations.
  • Formal et al. (2021) Thibault Formal, Carlos Lassance, Benjamin Piwowarski, and Stéphane Clinchant. 2021. Splade v2: Sparse lexical and expansion model for information retrieval. arXiv preprint arXiv:2109.10086.
  • Gao et al. (2021) Tianyu Gao, Xingcheng Yao, and Danqi Chen. 2021. SimCSE: Simple contrastive learning of sentence embeddings. In Proceedings of the 2021 Conference on Empirical Methods in Natural Language Processing, pages 6894–6910, Online and Punta Cana, Dominican Republic. Association for Computational Linguistics.
  • Gilardi et al. (2023) Fabrizio Gilardi, Meysam Alizadeh, and Maël Kubli. 2023. Chatgpt outperforms crowd-workers for text-annotation tasks. arXiv preprint arXiv:2303.15056.
  • Greene et al. (2022) Ryan Greene, Ted Sanders, Lilian Weng, and Arvind Neelakantan. 2022. New and improved embedding model. https://openai.com/blog/new-and-improved-embedding-model/. Accessed: 2023-06-03.
  • Hämäläinen et al. (2023) Perttu Hämäläinen, Mikke Tavast, and Anton Kunnari. 2023. Evaluating large language models in generating synthetic hci research data: a case study. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems, CHI ’23, New York, NY, USA. Association for Computing Machinery.
  • Izacard et al. (2022) Gautier Izacard, Mathilde Caron, Lucas Hosseini, Sebastian Riedel, Piotr Bojanowski, Armand Joulin, and Edouard Grave. 2022. Unsupervised dense information retrieval with contrastive learning. Transactions on Machine Learning Research.
  • Jeronymo et al. (2023) Vitor Jeronymo, Luiz Bonifacio, Hugo Abonizio, Marzieh Fadaee, Roberto Lotufo, Jakub Zavrel, and Rodrigo Nogueira. 2023. Inpars-v2: Large language models as efficient dataset generators for information retrieval. arXiv preprint arXiv:2301.01820.
  • Josifoski et al. (2023) Martin Josifoski, Marija Sakota, Maxime Peyrard, and Robert West. 2023. Exploiting asymmetry for synthetic training data generation: Synthie and the case of information extraction. arXiv preprint arXiv:2303.04132.
  • Kojima et al. (2022) Takeshi Kojima, Shixiang (Shane) Gu, Machel Reid, Yutaka Matsuo, and Yusuke Iwasawa. 2022. Large language models are zero-shot reasoners. In Advances in Neural Information Processing Systems, volume 35, pages 22199–22213. Curran Associates, Inc.
  • Li et al. (2023) Zhuoyan Li, Hangxiao Zhu, Zhuoran Lu, and Ming Yin. 2023. Synthetic data generation with large language models for text classification: Potential and limitations. arXiv preprint arXiv:2310.07849.
  • Liang et al. (2020) Davis Liang, Peng Xu, Siamak Shakeri, Cicero Nogueira dos Santos, Ramesh Nallapati, Zhiheng Huang, and Bing Xiang. 2020. Embedding-based zero-shot retrieval through query generation. arXiv preprint arXiv:2009.10270.
  • Lin et al. (2023) Kevin Lin, Kyle Lo, Joseph E Gonzalez, and Dan Klein. 2023. Decomposing complex queries for tip-of-the-tongue retrieval. arXiv whapreprint arXiv:2305.15053.
  • Liu et al. (2023) Jiduan Liu, Jiahao Liu, Qifan Wang, Jingang Wang, Wei Wu, Yunsen Xian, Dongyan Zhao, Kai Chen, and Rui Yan. 2023. RankCSE: Unsupervised sentence representations learning via learning to rank. In Proceedings of the 61st Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), pages 13785–13802, Toronto, Canada. Association for Computational Linguistics.
  • Liu et al. (2019) Yinhan Liu, Myle Ott, Naman Goyal, Jingfei Du, Mandar Joshi, Danqi Chen, Omer Levy, Mike Lewis, Luke Zettlemoyer, and Veselin Stoyanov. 2019. Roberta: A robustly optimized bert pretraining approach. arXiv preprint arXiv:1907.11692.
  • Loshchilov and Hutter (2017) Ilya Loshchilov and Frank Hutter. 2017. Decoupled weight decay regularization. arXiv preprint arXiv:1711.05101.
  • Ma et al. (2021) Ji Ma, Ivan Korotkov, Yinfei Yang, Keith Hall, and Ryan McDonald. 2021. Zero-shot neural passage retrieval via domain-targeted synthetic question generation. In Proceedings of the 16th Conference of the European Chapter of the Association for Computational Linguistics: Main Volume, pages 1075–1088, Online. Association for Computational Linguistics.
  • Meng et al. (2022) Rui Meng, Ye Liu, Semih Yavuz, Divyansh Agarwal, Lifu Tu, Ning Yu, Jianguo Zhang, Meghana Bhat, and Yingbo Zhou. 2022. Augtriever: Unsupervised dense retrieval by scalable data augmentation. arXiv preprint arXiv:2212.08841.
  • Nguyen et al. (2016) Tri Nguyen, Mir Rosenberg, Xia Song, Jianfeng Gao, Saurabh Tiwary, Rangan Majumder, and Li Deng. 2016. Ms marco: A human generated machine reading comprehension dataset. choice, 2640:660.
  • Peng et al. (2023) Zhiyuan Peng, Xuyang Wu, and Yi Fang. 2023. Soft prompt tuning for augmenting dense retrieval with large language models. arXiv preprint arXiv:2307.08303.
  • Penha et al. (2023) Gustavo Penha, Enrico Palumbo, Maryam Aziz, Alice Wang, and Hugues Bouchard. 2023. Improving content retrievability in search with controllable query generation. In Proceedings of the ACM Web Conference 2023, WWW ’23, page 3182–3192, New York, NY, USA. Association for Computing Machinery.
  • Robertson et al. (2009) Stephen Robertson, Hugo Zaragoza, et al. 2009. The probabilistic relevance framework: Bm25 and beyond. Foundations and Trends® in Information Retrieval, 3(4):333–389.
  • Santhanam et al. (2022) Keshav Santhanam, Omar Khattab, Jon Saad-Falcon, Christopher Potts, and Matei Zaharia. 2022. Colbertv2: Effective and efficient retrieval via lightweight late interaction. In Proceedings of the 2022 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, pages 3715–3734.
  • Sparck Jones (1972) Karen Sparck Jones. 1972. A statistical interpretation of term specificity and its application in retrieval. Journal of documentation, 28(1):11–21.
  • Sun et al. (2023) Weiwei Sun, Lingyong Yan, Xinyu Ma, Pengjie Ren, Dawei Yin, and Zhaochun Ren. 2023. Is chatgpt good at search? investigating large language models as re-ranking agent. arXiv preprint arXiv:2304.09542.
  • Thakur et al. (2021) Nandan Thakur, Nils Reimers, Andreas Rücklé, Abhishek Srivastava, and Iryna Gurevych. 2021. Beir: A heterogeneous benchmark for zero-shot evaluation of information retrieval models. In Thirty-fifth Conference on Neural Information Processing Systems Datasets and Benchmarks Track (Round 2).
  • Veselovsky et al. (2023) Veniamin Veselovsky, Manoel Horta Ribeiro, Akhil Arora, Martin Josifoski, Ashton Anderson, and Robert West. 2023. Generating faithful synthetic data with large language models: A case study in computational social science. arXiv preprint arXiv:2305.15041.
  • Wachsmuth et al. (2018) Henning Wachsmuth, Shahbaz Syed, and Benno Stein. 2018. Retrieval of the best counterargument without prior topic knowledge. In Proceedings of the 56th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), pages 241–251, Melbourne, Australia. Association for Computational Linguistics.
  • Wadden et al. (2020) David Wadden, Shanchuan Lin, Kyle Lo, Lucy Lu Wang, Madeleine van Zuylen, Arman Cohan, and Hannaneh Hajishirzi. 2020. Fact or fiction: Verifying scientific claims. In Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing (EMNLP), pages 7534–7550, Online. Association for Computational Linguistics.
  • Wang et al. (2023) Jianyou Wang, Kaicheng Wang, Xiaoyue Wang, Prudhviraj Naidu, Leon Bergen, and Ramamohan Paturi. 2023. Doris-mae: Scientific document retrieval using multi-level aspect-based queries. In Thirty-seventh Conference on Neural Information Processing Systems Datasets and Benchmarks Track.
  • Wang et al. (2022a) Kexin Wang, Nandan Thakur, Nils Reimers, and Iryna Gurevych. 2022a. GPL: Generative pseudo labeling for unsupervised domain adaptation of dense retrieval. In Proceedings of the 2022 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, pages 2345–2360, Seattle, United States. Association for Computational Linguistics.
  • Wang et al. (2022b) Liang Wang, Nan Yang, Xiaolong Huang, Binxing Jiao, Linjun Yang, Daxin Jiang, Rangan Majumder, and Furu Wei. 2022b. Simlm: Pre-training with representation bottleneck for dense passage retrieval. arXiv preprint arXiv:2207.02578.
  • Wang et al. (2022c) Liang Wang, Nan Yang, Xiaolong Huang, Binxing Jiao, Linjun Yang, Daxin Jiang, Rangan Majumder, and Furu Wei. 2022c. Text embeddings by weakly-supervised contrastive pre-training.
  • Wei et al. (2021) Jason Wei, Maarten Bosma, Vincent Zhao, Kelvin Guu, Adams Wei Yu, Brian Lester, Nan Du, Andrew M Dai, and Quoc V Le. 2021. Finetuned language models are zero-shot learners. In International Conference on Learning Representations.
  • Wei et al. (2022a) Jason Wei, Yi Tay, Rishi Bommasani, Colin Raffel, Barret Zoph, Sebastian Borgeaud, Dani Yogatama, Maarten Bosma, Denny Zhou, Donald Metzler, et al. 2022a. Emergent abilities of large language models. arXiv preprint arXiv:2206.07682.
  • Wei et al. (2022b) Jason Wei, Xuezhi Wang, Dale Schuurmans, Maarten Bosma, brian ichter, Fei Xia, Ed Chi, Quoc V Le, and Denny Zhou. 2022b. Chain-of-thought prompting elicits reasoning in large language models. In Advances in Neural Information Processing Systems, volume 35, pages 24824–24837. Curran Associates, Inc.
  • Weisstein (2004) Eric W Weisstein. 2004. Bonferroni correction. https://mathworld. wolfram. com/.
  • Xiong et al. (2020) Lee Xiong, Chenyan Xiong, Ye Li, Kwok-Fung Tang, Jialin Liu, Paul Bennett, Junaid Ahmed, and Arnold Overwijk. 2020. Approximate nearest neighbor negative contrastive learning for dense text retrieval. arXiv preprint arXiv:2007.00808.
  • Yang et al. (2023) Kevin Yang, Dan Klein, Asli Celikyilmaz, Nanyun Peng, and Yuandong Tian. 2023. Rlcd: Reinforcement learning from contrast distillation for language model alignment. arXiv preprint arXiv:2307.12950.
8.   Language Resource References
[…]
  • Agirre et al. (2013) Eneko Agirre and Daniel Cer and Mona Diab and Aitor Gonzalez-Agirre and Weiwei Guo. 2013. Semantic Textual Similarity (STS) 2013 Machine Translation. Linguistic Data Consortium, ISLRN 857-492-590-583-5. PID https://doi.org/10.35111/cy4d-7c39.
  • Kwiatkowski et al. (2019) Kwiatkowski, Tom and Palomaki, Jennimaria and Redfield, Olivia and Collins, Michael and Parikh, Ankur and Alberti, Chris and Epstein, Danielle and Polosukhin, Illia and Devlin, Jacob and Lee, Kenton and Toutanova, Kristina and Jones, Llion and Kelcey, Matthew and Chang, Ming-Wei and Dai, Andrew M. and Uszkoreit, Jakob and Le, Quoc and Petrov, Slav. 2019. Natural Questions: A Benchmark for Question Answering Research. Transactions of the Association for Computational Linguistics. PID https://doi.org/10.1162/tacl_a_00276.
  • Lin et al. (2023) Lin, Kevin and Lo, Kyle and Gonzalez, Joseph E and Klein, Dan. 2023. Decomposing Complex Queries for Tip-of-the-tongue Retrieval. arXiv preprint arXiv:2305.15053. PID https://huggingface.co/datasets/nlpkevinl/whatsthatbook.
  • Voorhees et al. (2020) Voorhees, Ellen and Alam, Tasmeer and Bedrick, Steven and Demner-Fushman, Dina and Hersh, William R. and Lo, Kyle and Roberts, Kirk and Soboroff, Ian and Wang, Lucy Lu. 2020. TREC-COVID: constructing a pandemic information retrieval test collection. ACM SIGIR Forum. PID https://doi.org/10.1145/3451964.3451965.
  • Wachsmuth et al. (2018) Henning Wachsmuth and Shahbaz Syed and Benno Stein. 2018. ArguAna Counterargs. Zenodo. PID https://doi.org/10.5281/zenodo.3973258.
  • Wang et al. (2023) Wang, Jianyou and Wang, Kaicheng and Wang, Xiaoyue and Naidu, Prudhviraj and Bergen, Leon and Paturi, Ramamohan. 2023. DORIS-MAE-v1. Zenodo. PID https://doi.org/10.5281/zenodo.8299749.

### v2

@@ 5.1 Results on WhatsThatBook
5.2 Results on DORIS-MAE
@@ 5.3 Results on ArguAna
Method R@5 R@20 NDCG@10 MRR@10 R@5 R@20 NDCG@10 MRR@10
Pretrained 59.37 88.16 47.75 38.96 30.51 49.68 23.56 18.49
Contrastive Fine-tuning 70.69 94.77 56.15 47.13 59.12 84.67 47.57 39.65
Promptagator 69.41 94.11 55.14 46.67 61.51 88.19 49.80 41.85
Doc40% reg 78.72 $\uparrow$ 9.31 96.46 $\uparrow$ 2.35 64.24 $\uparrow$ 9.10 56.45 $\uparrow$ 9.78 67.38 $\uparrow$ 5.87 91.03 $\uparrow$ 2.84 54.54 $\uparrow$ 4.74 46.43 $\uparrow$ 4.58
Queryreg $\circ$ Doc40% reg 77.76 $\uparrow$ 8.35 96.21 $\uparrow$ 2.10 63.00 $\uparrow$ 7.86 55.24 $\uparrow$ 8.57 67.29 $\uparrow$ 5.78 91.08 $\uparrow$ 2.89 54.04 $\uparrow$ 4.24 45.85 $\uparrow$ 4.00
Doc60% reg 78.16 $\uparrow$ 8.75 96.41 $\uparrow$ 2.30 64.41 $\uparrow$ 9.27 56.57 $\uparrow$ 9.90 67.51 $\uparrow$ 6.00 90.77 $\uparrow$ 2.58 54.85 $\uparrow$ 5.05 46.85 $\uparrow$ 5.00
Queryreg $\circ$ Doc60% reg 78.34 $\uparrow$ 8.93 96.37 $\uparrow$ 2.26 64.10 $\uparrow$ 8.96 56.24 $\uparrow$ 9.57 67.71 $\uparrow$ 6.20 90.99 $\uparrow$ 2.80 54.41 $\uparrow$ 4.61 46.34 $\uparrow$ 4.49
Doc80% reg 77.37 $\uparrow$ 7.96 95.57 $\uparrow$ 1.46 63.14 $\uparrow$ 8.00 55.16 $\uparrow$ 8.49 67.82 $\uparrow$ 6.31 90.18 $\uparrow$ 1.99 54.62 $\uparrow$ 4.82 46.79 $\uparrow$ 4.94
Queryreg $\circ$ Doc80% reg 78.44 $\uparrow$ 9.03 96.25 $\uparrow$ 2.14 64.36 $\uparrow$ 9.22 56.49 $\uparrow$ 9.82 67.38 $\uparrow$ 5.87 90.70 $\uparrow$ 2.51 54.53 $\uparrow$ 4.73 46.59 $\uparrow$ 4.74
Instrreg 72.06 $\uparrow$ 2.65 95.70 $\uparrow$ 1.59 57.76 $\uparrow$ 2.62 49.17 $\uparrow$ 2.50 64.47 $\uparrow$ 2.96 90.09 $\uparrow$ 1.90 51.49 $\uparrow$ 1.69 43.20 $\uparrow$ 1.35
Queryreg $\circ$ Instrreg 71.59 $\uparrow$ 2.18 95.18 $\uparrow$ 1.07 56.89 $\uparrow$ 1.75 48.37 $\uparrow$ 1.70 65.14 $\uparrow$ 3.63 90.26 $\uparrow$ 2.07 52.25 $\uparrow$ 2.45 43.89 $\uparrow$ 2.04
@@
Method R@5 R@20 NDCG@10 MRR@10 R@5 R@20 NDCG@10 MRR@10
Pretrained 49.03 81.12 39.23 30.35 38.91 71.79 30.91 23.23
Contrastive Fine-tuning 36.70 64.79 28.85 22.12 41.90 75.70 33.20 25.26
Promptagator 69.49 92.54 56.38 47.97 38.64 75.15 30.24 21.77
Doc40% reg 74.92 $\uparrow$ 5.43 93.36 $\uparrow$ 0.82 60.24 $\uparrow$ 3.86 52.42 $\uparrow$ 4.45 56.09 $\uparrow$ 17.45 86.32 $\uparrow$ 11.17 45.00 $\uparrow$ 14.76 36.30 $\uparrow$ 14.53
Queryreg $\circ$ Doc40% reg 76.07 $\uparrow$ 6.58 93.97 $\uparrow$ 1.43 61.38 $\uparrow$ 5.00 53.57 $\uparrow$ 5.60 51.28 $\uparrow$ 12.64 83.33 $\uparrow$ 8.18 40.59 $\uparrow$ 10.35 31.81 $\uparrow$ 10.04
Doc60% reg 73.20 $\uparrow$ 3.71 92.10 58.46 $\uparrow$ 2.08 50.58 $\uparrow$ 2.61 58.87 $\uparrow$ 20.23 86.93 $\uparrow$ 11.78 46.55 $\uparrow$ 16.31 37.88 $\uparrow$ 16.11
Queryreg $\circ$ Doc60% reg 75.75 $\uparrow$ 6.26 93.19 $\uparrow$ 0.65 60.94 $\uparrow$ 4.56 53.24 $\uparrow$ 5.27 54.22 $\uparrow$ 15.58 84.97 $\uparrow$ 9.82 42.90 $\uparrow$ 12.66 34.21 $\uparrow$ 12.44
Doc80% reg 70.38 89.67 $\downarrow$ -2.87 55.76 47.97 61.02 $\uparrow$ 22.38 87.48 $\uparrow$ 12.33 48.68 $\uparrow$ 18.44 40.24 $\uparrow$ 18.47
Queryreg $\circ$ Doc80% reg 74.10 $\uparrow$ 4.61 92.59 59.79 $\uparrow$ 3.41 52.18 $\uparrow$ 4.21 57.55 $\uparrow$ 18.91 86.65 $\uparrow$ 11.50 46.01 $\uparrow$ 15.77 37.44 $\uparrow$ 15.67
Instrreg 70.97 93.91 $\uparrow$ 1.37 57.41 48.80 43.26 $\uparrow$ 4.62 80.16 $\uparrow$ 5.01 34.80 $\uparrow$ 4.56 26.07 $\uparrow$ 4.30
Queryreg $\circ$ Instrreg 71.46 $\uparrow$ 1.97 93.78 $\uparrow$ 1.24 58.16 $\uparrow$ 1.78 49.78 $\uparrow$ 1.81 38.57 77.12 $\uparrow$ 1.97 31.72 $\uparrow$ 1.48 23.15 $\uparrow$ 1.38